OpenStudy (jekdidkdjrjrjjr):

FAN+MEDAL What is 120 degrees in radians?

OpenStudy (jekdidkdjrjrjjr):

options

OpenStudy (jekdidkdjrjrjjr):

@Atrineas @ScarlettFarra2000 @anthonyym @SamandDean1967

OpenStudy (cinnamonbun):

\[\frac{ 2\pi }{ 3 }\]

OpenStudy (jekdidkdjrjrjjr):

u sure?

OpenStudy (cinnamonbun):

Yep

OpenStudy (anthonyym):

There are 2pi radians in a circle, therefore 360 degrees = 2pi Or divide by 2 and get. 180 degrees = pi Knowing that 180 degrees = pi, this will be your conversion factor. Multiply 120 degrees by the conversion factor pi/180 (because pi/180 = 1) \[\frac{ 120 }{ 1 }*\frac{ \pi }{ 180 }\] So this is the same as (120/180)*pi

OpenStudy (anthonyym):

Which you can simplify to (3/4)pi